OpenStudy (anonymous):

hat is the value of the expression when a = 5 and b = 3? 4a^2 5b^3 A. 55 B. 74 C. 81 D. 145

OpenStudy (anonymous):

@jim_thompson5910 pls help

OpenStudy (anonymous):

@jim_thompson5910 pls help

jimthompson5910 (jim_thompson5910):

Replace 'a' in 4a^2*5b^3 with 5 so we have 4*5^2*5b^3

jimthompson5910 (jim_thompson5910):

also replace b with 3 to get 4*5^2*5*3^3 I'll let you finish

OpenStudy (anonymous):

okay 100-45=55?

jimthompson5910 (jim_thompson5910):

oh the original expression is \[\Large 4a^2 - 5b^3\] ??

OpenStudy (anonymous):

right

jimthompson5910 (jim_thompson5910):

I'm not sure how you got 45

jimthompson5910 (jim_thompson5910):

oh I see. You did 3*3 = 9 when you should have done 3^3 = 3*3*3 = 27

jimthompson5910 (jim_thompson5910):

try it again

OpenStudy (anonymous):

its 55 thanks tho